Find the product using suitable properties
258 x 1008

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given 258 × 1008
= 258 × (1000 + 8)
= 258 × 1000 + 258 × 8 (using distributive property)​
= 258000 + 2064
= 260064

Thanks